OpenStudy (anonymous):

The x co-ordinate of a point on the curve 9y^2 = x^3 , the normal at which cuts off equal intercept on the axes is ?

OpenStudy (anonymous):

@shubhamsrg

OpenStudy (anonymous):

dy/dx =x^2/(6y)

OpenStudy (anonymous):

gradient of normal is -6y/(x^2)

OpenStudy (anonymous):

Nw wat to do ?

OpenStudy (shubhamsrg):

-6y/x^2 is the slope of a line which cuts of equal intercepts as x axis.

OpenStudy (shubhamsrg):

hence slope should be equal to 1 or -1.

OpenStudy (shubhamsrg):

"-6y/x^2 is the slope of a line which cuts of equal intercepts ON THE axes.

OpenStudy (shubhamsrg):



OpenStudy (anonymous):

Did nt get u

OpenStudy (shubhamsrg):

-6y/x^2 is the slope of the normal right ?

OpenStudy (anonymous):

yup

OpenStudy (shubhamsrg):

and we are given that the normal cuts of equal intercepts on the axes.

OpenStudy (anonymous):

ok...then

OpenStudy (anonymous):

let the Points be (a,0) and (0,a ) ryt ?

OpenStudy (shubhamsrg):

any line which cuts off equal intercepts on the axes must have a slope of 1 or -1 right? Since it is making an angle of pi/4 or 3pi/4 with the +ve X axes.

OpenStudy (anonymous):

(a-0) / (0-a) = -1

OpenStudy (shubhamsrg):

What I mean is -6y /x^2 = 1 and/or -1

OpenStudy (anonymous):

ok...

OpenStudy (shubhamsrg):

we also know y = x^(3/2) /3 or -x^(3/2) /3

OpenStudy (anonymous):

yes

OpenStudy (shubhamsrg):

substitute in the previous eqn to find the required value(s) of x

OpenStudy (anonymous):

Yeah got it..:) thxx

OpenStudy (shubhamsrg):

glad you did. ^_^